hw 9 # 17

Data on the numbers of hospital admissions resulting from motor vehicle crashes are given below for Fridays on the 6th of a month and Fridays on the following 13th of the same month. Assume that the paired sample data is a simple random sample and that the differences have a distribution that is approximately normal. Construct a​ 95% confidence interval estimate of the mean of the population of differences between hospital admissions. Use the confidence interval to test the claim that when the 13th day of a month falls on a​ Friday, the numbers of hospital admissions from motor vehicle crashes are not affected.

Friday the 6th

2

6

6

4

8

Friday the 13 thFriday the 13th

14

15

11

14

15

In this​ example,

mu Subscript dμd

is the mean value of the differences d for the population of all pairs of​ data, where each individual difference d is defined as the number of hospital admissions on Friday the 6th minus the number of hospital admissions on Friday the 13th. Find the​ 95% confidence interval.
Answer #1

Ans:

Friday the 6th Friday the 13th d
2 14 -12
6 15 -9
6 11 -5
4 14 -10
8 15 -7
d-bar -8.6
sd 2.70185

df=5-1=4

critical t value=tinv(0.05,4)=2.7764

Margin of error=2.7764*(2.70185/SQRT(5))=3.35

95% confidence interval for mean difference

=-8.6+/-3.35

=(-11.95, -5.25)